OpenStudy (anonymous):

can someone please help me solve this limit lim [1/(8+x) -1/8]/x X-->0

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{\frac{1}{8+x}-\frac{1}{8}}{x}\] It is a \[\frac{0}{0}\] You probably should apply L'Hopital \[\lim_{x \rightarrow 0} \frac{\frac{1}{-(8+x)^2}-0}{1}=\lim_{x \rightarrow 0} -\frac{1}{(8+x)^2}=\frac{1}{64}\]

OpenStudy (anonymous):

have never seen the law , i was trying to get common denom and cancel out but got stuck, thanks still